A SEMIEXPERIMENTAL EQUILIBRIUM STRUCTURE OF cis-HEXATRIENE FROM MICROWAVE SPECTROSCOPY

\underline{NORMAN C. CRAIG}, YIHUI CHEN, HANNAH A. FUSON, HENGFENG TIAN, and HERMAN VAN BESIEN, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, Ohio 44074; H. D. RUDOLPH, Department of Chemistry, University of Ulm, D-89069 Ulm, Germany; JEAN DEMAISON, Laboratoire de Physique des Lasers, Atomes et Molécules, Université de Lille I, 59655 Villeneuve d’Ascq Cedex, France.

Ground state rotational constants for a full set of carbon and deuterium isotopologues of cis-hexatriene, which has a dipole moment of only 0.05 D, have been determined by MW spectroscopy.^{a}^{b} Ground state rotational constants were converted into equilibrium rotational constants with spectroscopic alphas computed from force constants found with an MP2/ccpVTZ model. Corrections for electronic contributions were applied. Despite the availability of a full set of rotational constants for a suite of isotopologues, an accurate structure was not accessible. The mixed estimation method, in which high-quality bond parameters computed by quantum chemistry with appropriate uncertainties were fit simultaneously with the moments of inertia, was applied. A structure good to 0.001 Å was the result. The structural parameters reflect an increase in pi-electron delocalization compared to butadiene.

aR. D. Suenram, B. H. Pate, A. Lesarri, J. L. Neill, S. Shipman, R. A. Holmes, M. C. Leyden, N. C. Craig J. Phys. Chem. A 113, 1864-1868 (2009).

bN. C. Craig, Y. Chen, H. A. Fuson, H. Tian, H. van Besien, A. R. Conrad, M. J. Tubergen, H. D. Rudolph, J. Demaison, J.

Phys. Chem. A 117, in press.